UT Feb 18 2022 - Split Night - LBTI 1st half - DEC 2nd half



Staff

  • OSA: S. Allanson
  • ISp: C.Bolyard
  • Mtn Manager: B.Bingham
  • ISA/Observer: J.Power
  • AO: G.Taylor
  • Special Guests: B.Rothberg, R.T. Gatto

Night Info

  • Sunset 18:07
  • Nautical Twilight Ends 19:01
  • Moon rise 19:13
  • Astro Twilight Ends 19:30
  • Night Center: 00:33 - Handover from LBTI to LUCI-AO (LUCI-Bino) 00:27
  • Astro Twilight Starts 05:37
  • Nautical Twilight Start 06:06
  • Sunrise 06:59
  • Moon illumination 96.8%

Night Plan

LUCI-1 AO/ESM

  • Install ao-supervisor "latest" version
    • see B.Mechtley's email notes on installation and rollback
  • OSCO - 00:30-01hr
  • General LUCI sub-frame test - change filters via a script when using the RTD to sub-frame
  • Test the ramp/resume of the new software - 0.75hr
    • Use any AO on-axis test star - make sure to run with offsets lasting at least 1 minute of shutter time (Bin 3 or BIn 4 are best)
    • Test with 2 different stars to verify gains ramp up correctly on return
  • Bin 1 star with NCPA on/off at Instrument Rotator ~360deg - 0.5hr
    • UT Feb 14 found better IQ with Bin 1 NCPA off at rotator~360deg - repeat test to confirm
      • Calculate and select possible targets during night - want something around 360
        • ROTANGLE = PA - PARANGLE + (180 - LEFTZEROPOINT)
        • Left ZP =-187.735
      • Quick: Rotangle = PA - PARANGLE + 7.735
  • Moon Angle test -1hr
    • 10 degrees from Moon
      • Bin 1 or Bin 2 10 deg from Moon
        • RA 10-11hr folder Star #9 (2MASS J11010240-0030040 M-star RGB) R=10.42 (11.4)
      • Bin 3 or Bin 4 10 deg from Moon
        • RA 10h-11hr folder, Star #4 PA 240 (GD 135 WD) R=14.6 (15.8)
    • 15 deg from Moon
      • Bin 3 or Bin 4
        • RA 10-11hr folder, Star #10 PA 180 (G 58-40) R=13.9 (15.1)
  • ESM
    • Verify updated TTM table works for BIn 2
      • M67 R=12.2 (change to R=12.6 on script) - visible handover-03:30
      • Dual BLR SDSSJ1713 - R=12.3 (change to R=12.6 on script) - visible 03hr-twilight
    • Bin 1 target
      • Verify no re-rotator or OCAM filter wheel problems
      • M61 - face-on spiral R=10.47 star visible handover-twilight
  • AO "Pretty Pix"
    • Cat's Eye Nebula - On axis
      • R=11.38 (12.58 - Bin 1 or Bin 2 - borderlien depending on WFSmag)
      • Visible 03hr-twilight
      • Ks, H2, Pbeta, FeII
  • Roll back ao-supervisor "latest" version after closing up

Backup

  • Poor Weather
    • Local Mergers - Rothberg
      • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V40.xml
      • Any conditions (guiders up to 4" - as long as guider can lock on and WFS) - no particular order
        • UGC 9829 - 2nd half
        • UGC 10607 -2nd half
        • UGC 10675 -2nd half
        • NGC 4004 -2nd half
        • Arp 156 -2nd half
        • NGC 3597 -00:00-03:30hr
      • If clear and super poor-poor seeing
        • LUCI-Bino Standards
          • LUCI - /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards/LUCI_Phot_ZP_OT_V3.xml
          • If seeing > 4" - make these track modes in the OT and export
  • If AO cannot be execute and conditions are good
    • LUCI-Bino SL - GRB 211211A - 1.25hr (2nd half) - Rothberg
      • OT File: /home/lbto/Service_Observing/2022A_D_E/LUCI/GRB211211A/GRB211211A_LD2021B-002_V5_OT.xml
      • Ks and J (fraternal) imaging
      • Seeing < 1.3"
      • Thin Clouds
      • If clear (photometric) - execute standard GSPC P272-D before and after observations (just rotates through Ks and J)

Night Log

7:27 Handover from LBTI. Steve moving to zenith out of the wind so we can restart the adsec with the new software

7:44 The adsec arbitrator is not coming up. Brandon is on working his magic.

Brandon fixed the issues. the arbitrator came up ok, the new force open loop button is

8:00 Steve pointing and collimating after the new software has been deployed successfully. Seeing ~1.5". Conditions clear.

8:09 Slewing to OSCO target HD71514 R 7.52, elevation49, rotator 141 to start. We did not start collimating right away. It claimed it was moving the probe after the fact, probe position may have drifted, but it recovered after ~30 sec. Seeing 1.2" on the guider.

WFS mag 8.7
Fri Feb 18 08:10:40.402 2022 59628.34074539813 2 GCSL gcs.guidethread.alert left Probe position difference between requested/achieved is large 0.517379 (arcs). Probe position may have DRIFTED which may cause WFS to pause. 2022-02-18 08:13:21.275124| -13006 - [AOException] centerPupils: WARNING - RETRY: Camera lens temperature too low. T= 6.89, minimum allowed is 7.00 (0) NO_ERROR

Steve is running the FLAOGUI to close the loop.

Center pupils failed the first time
WfsArbitrator_98663|ERR| 144|2022-02-18 08:13:15.868706| MAIN > RETRY: Camera lens temperature too low. T= 6.89, minimum allowed is 7.00 (0) NO_ERROR

Repeat successful.

Optimize gain routine failed first time:
WfsArbitrator_98663|ERR| 212|2022-02-18 08:15:53.023883| MAIN > Data analysis routine (in IDL) did not finish correctly (-3006) VAR_NOT_FOUND_ERROR

luci1.20220218.0063-82 NCPA off
NO NCPA, 1st POSITION:
=> Command line arguements processed
Set 0 Tracknum 20220218_082202 LuciName unknown
Set 1 Tracknum 20220218_082215 LuciName unknown
Set 2 Tracknum 20220218_082226 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220218_082202.log

NO NCPA, 2nd POSITION:
=> Command line arguements processed
Set 0 Tracknum 20220218_082444 LuciName unknown
Set 1 Tracknum 20220218_082457 LuciName unknown
Set 2 Tracknum 20220218_082510 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220218_082444.log

The rolled edge appeared as we turned on NCPA. High number of skip frames. Review Optical loop data to see why so many skip frames, likely high forces.

luci1.20220218.0083-0101 NCPA on
NCPA ON, 1st POSITION (some skip frames):
=> Command line arguements processed
Set 0 Tracknum 20220218_083028 LuciName unknown
Set 1 Tracknum 20220218_083041 LuciName unknown
Set 2 Tracknum 20220218_083053 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220218_083028.log

NCPA ON, 2nd POSITION (some skip frames):
=> Command line arguements processed
Set 0 Tracknum 20220218_083428 LuciName unknown
Set 1 Tracknum 20220218_083440 LuciName unknown
Set 2 Tracknum 20220218_083452 LuciName unknown
-> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220218_083428.log

8:40 Slewing to star1 in RA14-15, bin1 PA60. Concerns about wind so abroting and slewing to next candidate

8:43 Slewing to Star 5 in in RA14-15, bin1 PA60. No guide star first time, so regenerated and resent. We have a glint from the moon that is causing hte wfs to be filled.

8:51 Executing Moon test instead. Star 9 in RA10-11 folder R=10.4. We will be 10.9 degrees from the moon. The guide star is too faint and we are out of range for the rerotator. Selecting a new guide star at PA270

8:58. New guide star of 12.25 clearly visible. This was needed since we fell out of collimation. Rotator now at 265 and decreasing, seeing 1.3-1.7"

R mag 10.4, OCAM seeing 11.2

Rolled edge seen when turning optical gain on (before NCPA on). Conditions are not very stable.

luci1.20220218.0103 was accidentally full frame but we do see the jump in the integrated RTD image

We opened loop,. Unstable conditions by actuator 374 may be a detatched magnet.  
Actuator 374 appears to have detached magnet:
 Actuator374.png

9:14 Greg is removing actuator 374 from the list. adsec stop/start. Flat with 481 modes. That edge looks much better.

9:24 Resending preset to 2MASS J11010240240-0030040 for moon test, 10.9 degrees from moon. Seeing 1.2". R10.4, OCAM mag Rotator 237

massive tilt applied when centering pupils.

pupils also off when centering pupil so struggled to close loop. Greg had to manually adjust camera camera lens to align pupils. Suspect this occurred because background was so high OCAM mag 11.2.

Optimize gain drove way into safe skip and opened. The ranges are too big

MAJOR seeing bubble, highly variable

9:37 RIP from huge tilt during optimize gain.

9:44 Preset resent. Seeing variable. Can't keep loop closed

Executed subframed image with FeII 105 and BrGamma immediately after 106. Subframe did not reset.

9:56 Conditions are too unstable for AO. We are reauthorizing binocularly for full AO

9:58 Slewing to M67 ESM/67 forced bin2. Script magnitude R=12.6. Seeing 1.4-1.7. OCAM magnitude

Read error on LUCI2. Geirs was up not up. Conditions too unstable to keep loop closed

Greg verified the nifty new button on the flao gui works!

Moving on to poor weather backups

10:20 Slewing to APR second half telluric. Too low. LUCI2 mos error. Instrument Rotator moving to 210, el 90.
Grabber open, "rotator" at 340.977, prindle at 275.371, -4778

MOS storage to turnout, post condition violation translator to mask failed.

Check precon for init translator ok, init translator ok. Resolved and back in buisiness.

10:28 Slewing to Apr156 Telluric first half - HD95605. This target points us directly into the wind but winds ~2m/s on teh adsec, ~9/5m/s on the roof anemometer. Monitoring.

10:36 The guide star is way to bright and the guider is trying to guide on noise. Switched guide stars (PA90) and resending.

We are now getting 0.8-1.1" seeing on the guiders.

Science on HD95605

luci1.20220218.0110-114
luci2.20220218.0006-0008

10:50 Slewing to science field Arp156. Seing a bit variable but is about 1". Acquisition images taken and seeing is 0.8-1.1" at 40 degrees elevation. We are switching to AO

10:58 Steve is reauthorizing monocular luci

11:01 Slewing to bin3 star 4 R=14.6 PA=0 - GD135. The moon distance 10.5 degrees. Safe skipping. Center star took two tries. OCAM magnitude 15.5. Bin3, freq 754,

Optimize gain repeatedly fails. Gain plots are completely flat. Moonlight too high

11:12 Slewing to bin3 R13.9 Star 10 in RA10, PA180 - G58-40. Initial setup is setup as bin2, not bin3. Bin2, freq 836. OCAM magnitude. No issues closing the loop or optimizing, applying optical gain

banding in image 0116. This was a short 1x2.51 acquisition image. Images after this were 26x2.51 cubes. Field stop appears to have shifted. Will realign before we start next science target.
LUCI banding.png

Images 117-118 NCPA was not on. No banding seen. Turned on NCPA

We do not see any obvious sign gains are ramping after large offsets.

TN#114753 taken during offset.

12:01 Huge seeing bubble. Loop opened, but we are on teh last image of the sequence. Image 0131

12:05. Sending preset to cats eye nebula this is pointed intot he wind. Winds ~2 m/s at the adsec. Pointing check needed

12:07 Resending preset to Cats Eye nebula. Seeing 1.25" at 45 in elevation.

OCAM magnitude is 13.2. R mag is 11.38

Seeing variable while closing the loop. Up to 2". Barry jinxed it.

12:20 Shell rip while trying to optimize. Steve is slewing us to zenith to recover. Greg is going to revert the software.

12:27. Seeing too variable. Steve has reauthorized us for binocular luci and we are slewing to Arp156 science.

12:31 At 37 elevation the seeing is varying from 1.1-1.7"

luci1 acquisition images 0012 & 0013 did not appear in readout 2 or in the rrtd, but 0014 (slit image) sdid. The images were retrieved from newdata. Running on obs5. I am running the standard rtd. Barry noted the same feature, also running on obs5 but with the luci_rtd_test verion for luci2

13:08 Seeing has been holding around 0.8"
luci1.20220218.0142-147, luci2.20220218.0016-21

13:20 Closing up

Calibrations

G210 k@2.3 with 0.5" longslit luci1.20220218.0148-, luci2.20220218.0022-

MODS calibrations for INAF:

2021/MODS/Biazzo_13/cal

mos_592556_arc_flat_MODS1.txt and mos_592556_arc_flat_MODS2.txt

mods1b/r.20220218.0003-11
mods2b/r.20220218.0003-8, 0010-11
mods2b/r.20220218.0016 is Ne only

mods2[r/b].20220218.009Ne lamp failed to turn on, repeat: 0012-15 also failed, 0016 is turned on but is Ne only, ran script which cycled Ne lamp power and 0017 failed. 0018-25 tried Ne only again but failed.

IT #7869 priority raised. Unable to get the Ne lamp on.

-- BarryRothberg - 16 Feb 2022

  • Banding between channels in luci1 image:

*
Topic revision: r8 - 18 Feb 2022, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback